### v1

# Exploiting Meta-Learning-based Poisoning Attacks for Graph Link Prediction

Abstract
Link prediction in graph data utilizes various algorithms and machine learning/deep learning models to predict potential relationships between graph nodes. This technique has found widespread use in numerous real-world applications, including recommendation systems, community networks, and biological structures. However, recent research has highlighted the vulnerability of link prediction models to adversarial attacks, such as poisoning and evasion attacks. Addressing the vulnerability of these models is crucial to ensure stable and robust performance in link prediction applications. While many works have focused on enhancing the robustness of the Graph Convolution Network (GCN) model, the Variational Graph Auto-Encoder (VGAE), a sophisticated model for link prediction, has not been thoroughly investigated in the context of graph adversarial attacks. To bridge this gap, this article proposes an unweighted graph poisoning attack approach using meta-learning techniques to undermine VGAE’s link prediction performance. We conducted comprehensive experiments on diverse datasets to evaluate the proposed method and its parameters, comparing it with existing approaches in similar settings. Our results demonstrate that our approach significantly diminishes link prediction performance and outperforms other state-of-the-art methods.
{IEEEImpStatement}
Link prediction is a popular and critical technology in many applications, it helps organizations and researchers to understand and analyze the graph network. However, recent researches have pointed out that the link prediction model is vulnerable under adversarial attacks. Addressing the vulnerability of link prediction model can enhance the robustness of the link prediction model. The adversarial attack proposed in this paper reveals the vulnerabilities of a popular link prediction model (VGAE). With a minor and unnoticeable modification on the graph structure, the proposed adversarial approach can successfully degrade the performance of link prediction model. This approach indicates the possible threat to link prediction applications. It could provide a direction for enhancing the robustness of link prediction models.
{IEEEkeywords}
Adversarial attack, Link prediction, Meta-learning, Variational Graph Auto-encoder (VGAE)
1 Introduction
Many real-world datasets can be represented as graphs containing nodes, links, and associated features. Examples include social networks [1], traffic networks [2], biological structures [3], e-commerce networks [4], and so on. The structure of these networks is dynamic, changing over time as relationships evolve. Therefore, capturing these changes in the underlying graph is crucial for any application. Link prediction addresses this challenge and finds applications in various real-world scenarios, such as recommendation systems [5]. It employs diverse algorithms and machine learning (ML)/deep learning (DL) models to predict potential relationships (links) between nodes in a graph.
Over the past few decades, numerous approaches to link prediction have emerged across various research domains, including Local Random Walk [6], DeepWalk [7], node2vec [8], Graph Convolutional Networks (GCN) [9], Variational Graph Auto-encoder (VGAE) [10], and others. However, recent studies have underscored the susceptibility of graph learning models to adversarial attacks [11], such as poisoning and evasion attacks, which pose significant threats to the performance and stability of link prediction models. Given the widespread use of link prediction in real-world applications, ensuring the robustness and stability of these algorithms is essential. In essence, addressing the vulnerability of link prediction models is crucial, and this often falls within the realm of graph adversarial attacks.
In recent years, numerous approaches to graph adversarial attacks have been proposed. Xu et al. [12] proposed a projected gradient descent (PGD) topology attack from an optimization perspective to optimize the negative cross-entropy using the gradient to maximize the prediction error of the model. Waniek et al. [13] proposed the Disconnect Internally, Connect Externally (DICE) algorithm, which reduces internal graph density to deceive graph models. Chen et al. [14] leveraged an adversarial network generator to initiate attacks based on gradients, effectively misleading state-of-the-art link prediction algorithms. Furthermore, Meta-learning [15], offering valuable insights for efficient learning and rapid adaptation, has emerged as an attack strategy. In [16], the authors utilized meta-learning to conduct adversarial modifications on graph data.
Although minor alterations to graph data can significantly affect model performance, crafting an effective graph adversarial attack remains challenging. Attackers must meticulously design their strategies to ensure effectiveness. Additionally, as models and learning techniques advance, modern deep learning models exhibit increasing robustness, surpassing the resilience of previous generation machine learning models. For example, in [14], the author highlights the superior robustness of Graph Auto Encoder (GAE) compared to other state-of-the-art link prediction models. In addition, Variational Graph Auto-encoder (VGAE), which can be considered as an advancement of GAE has shown competitive performance on link prediction tasks. It uses a probabilistic approach, allowing the model to learn more flexible and interpretable latent representations of graph data. Nevertheless, the adversarial attack on VGAE remains a significant and open challenge.
In this study, we aim to bridge this gap by investigating adversarial attacks on VGAE and propose a solution through the introduction of an unweighted graph poisoning attack approach. Leveraging meta-learning techniques during the adversarial training stage, our method aims to degrade the link prediction performance of VGAE. Through extensive experiments on diverse datasets, we evaluated our approach and its parameters. In addition, we compare our method with existing approaches, demonstrating its superior effectiveness. To ensure reproducibility and facilitate further research, we have made the source code publicly available at our github repository [^1] https://github.com/mingchenli/VGAE_Attack_Meta.
The main contributions of our approach are summarized as follows:
  • We introduce an innovative poisoning attack leveraging meta-learning to effectively degrade the link prediction performance of VGAE.
  • By leveraging meta-learning, we determine optimal attack directions resulting in subtle modifications in the graph, while defining the adversarial behavior of these modifications.
  • Extensive experiments were conducted to assess the impact of parameters and the effectiveness of our approach across three benchmark datasets under various scenarios. Furthermore, we compare our approach with the state-of-the-art approaches to gauge its impact.
The remaining sections of the paper are structured as follows: Section 2 provides an overview of related works, Section 3 offers background knowledge on our approach, Section 4 outlines our methodology, and Section 5 presents the experimental evaluation, and Section 6 concludes the paper.
2 Related Works
Adversarial attacks on graph data have gained significant traction as a crucial topic in machine learning research in recent years. These attacks serve to identify model weaknesses and bolster model robustness through the development of corresponding defense mechanisms. Even minor and unnoticeable adjustments to graph data can lead to notable changes in model performance, underscoring the importance of addressing vulnerabilities in these machine learning models. Over the past few years, several effective attack approaches have been devised specifically for graph data and models.
2.1 Gradient-Based Attacks
Gradient-based adversarial attacks utilize model gradients to identify and perturb critical graph elements. Xu et al. [12] proposed projected gradient descent (PGD) and MinMax attacks that treat the attack as an optimization problem. They have proposed an optimization framework to optimize the negative cross-entropy using the gradient to maximize the prediction error of the model. Their method uses projected gradient descent (PGD) to iteratively modify links, significantly affecting model performance while respecting a budget constraint on the number of modifications. Chen et al. [14] introduced the iterative gradient attack (IGA) for the attack on link prediction tasks. This method leverages gradients derived from trained graph auto-encoders to identify and perturb the significant links in the graph. They perform comprehensive evaluations to reveal that even minor link alterations could disrupt the performance of state-of-the-art link prediction models. Zhang et al. [17] developed a gradient-based attack based on graph contrastive learning. By maximizing the contrastive loss during the training phase, the attack disrupts the embeddings produced by the contrastive model under an unsupervised learning setting, making it suitable for real-world where sometimes the data labels are scarce. In [16], a meta-learning based poisoning attack strategy was proposed to target the significant part of the graph structure or node characteristics. They using meta-gradient as guidance to generate adversarial modifications to decrease graph model performance.
2.2 Heuristic-Based Attacks
Heuristic-based attacks rely on predefined strategies to perturb graphs, often utilizing insights from graph theory. Zügner et al. proposed Nettack  [18], a framework that uses a greedy algorithm to identify and perturb key links and node attributes, forcing misclassifications in node classification tasks. Nettack shows that the weaknesses of the models are highly dependent on graph structure in semi-supervised settings. DICE attack introduced in [13], uses the strategies Disconnect Internally, Connect Externally (DICE) to manipulate the internal density within communities and disrupt specific targets in the graph. The attack modifies the graph structure strategically to jeopardize the performance of the graph model.
2.3 Reinforcement Learning-Based Attacks
Reinforcement learning-based methods utilize a sequential decision-making process with the agent to achieve the adversarial attack by using a reward system to maximize misclassification rates in black-box settings. Dai et al. [19] presented a reinforcement learning framework that modifies graph structures through link additions or deletions. Sun et al. [20] extended reinforcement learning to node injection attacks, where fake nodes are introduced into the graph. A reinforcement learning agent is used to determine the optimal connections for these nodes, maximizing the overall classification error while keeping the original structure of the graph similar.
However, most of these attacks mentioned above are targeted or designed for graph convolutional networks (GCNs) [9], which serve as the foundational model for graph data analysis. GCN has certain limitations on its representation capacity, graph data types, and generative capabilities. In contrast, VGAE has become a popular and more powerful model than GCNs for the link prediction task. Our approach presents a distinct meta-learning framework and proposes predefined attack modifications tailored for adversarial attacks in link prediction.
3 Background
While the detailed approach is presented in a later section, this section provides the relevant background knowledge for this paper.
3.1 Graph Adversarial Attack
In general, graph adversarial attacks utilize intentional modifications or perturbations on the graph structure, node attributes, or relevant features to degrade the performance of the target model. For graph data, these common modifications can be applied to the node, links (edges), and the whole graph, such as adding or deleting nodes of links, changing link weights (in weighted graph), modifying node features, and generating fake graphs.
Figure 1: Adversarial attacks based on the target stage of the ML pipeline.
Graph adversarial attacks can target machine learning/deep learning models at various stages throughout the learning pipeline. Based on the target stage, these attacks can be broadly categorized into two main categories (as shown in Fig. 1):
  • Evasion attack: This occurs during the model’s testing (inference) phase. The attacker manipulates the input data during testing to induce the model to make incorrect predictions. Adversarial examples are typically generated to test the vulnerability of the target model.
  • Poisoning attack: This occurs during the model’s training phase. The attacker injects malicious data into the original training dataset, producing poisoned data that bias the model’s learning process. The result is a dataset contaminated with malicious intent, used to train the target model.
In the realm of graph adversarial attacks, the objectives pursued by attackers are pivotal in shaping their strategies. These adversarial attacks can be further classified depending on the attacker’s goals and the context of the attack:
  • Untargeted attack: The objective is to reduce the overall prediction accuracy of the model in all instances.
  • Targeted attack: The objective is to specifically reduce the prediction accuracy of the model for certain instance(s).
Furthermore, based on the consequences of the attack, graph adversarial attacks can be classified into two main categories:
  • Node classification attack: This type of attack aims to manipulate the predictions of the machine learning/deep learning model, leading it to misclassify node(s) into incorrect class(es). For example, it might force a node in one community to be classified into another community within a social network graph.
  • Link prediction attack: The objective of this attack is to manipulate the model’s predictions regarding the likelihood of links between nodes. For example, the attack might cause the model to predict a relationship between two unrelated users in a social network graph, contrary to the actual scenario.
In this paper, our goal is to design a poisoning attack on link prediction tasks for both untargeted and target scenarios.
3.2 Variational Graph Auto-Encoders (VGAE)
The variational graph auto-encoder proposed by Kipf et al. [10] is a attractive choice for link prediction tasks due to its capability to learn meaningful latent representations for graph data and to directly reconstruct the graph structure via the decoder. VGAE’s objective is to maximize the Evidence Lower Bound (ELBO) $\mathcal{L}$ to enhance reconstruction accuracy while also regularizing the distribution of the latent space. Given the adjacency matrix $A$ and feature matrix $X$, the objective function of VGAE, derived from ELBO, can be expressed as:
$\mathcal{L}=\mathbb{E}_{q(Z|X,A)}[\log p(A|Z)]-KL(q(Z|X,A)||p(Z))$ (1)
where $\mathbb{E}$ denotes the expected value and $KL$ represents the Kullback-Leibler divergence [21] between the posterior $q$ and prior distribution $p$. KL divergence term acts as a regularizer, ensuring that the learned latent space is close to the prior distribution, which promotes better generalization and meaningful sampling.
In summary, the Variational Graph Autoencoder (VGAE) is a powerful generative model that learns to encode data into a probabilistic latent space and decode from this space to reconstruct the graph to perform link prediction or node classification tasks. Using KL divergence as a regularization term aims to balance reconstruction accuracy while ensuring that the latent space has known distribution and meaningful sampling. Due to these characteristics, the VGAE often outperforms conventional GCN-based approaches for link prediction, as it better captures the underlying distribution of the graph data. Hence, it is worth investigating the adversarial attack on VGAE.
3.3 Adversarial Attack with Meta-Learning
Meta-learning, also referred to as “learn to learn” [15], [22] is a widely used technology in the field of machine learning, recognized for its ability to facilitate efficient learning and rapid adaptation. Using knowledge acquired from various tasks, known as episodes, provides valuable insight into the model, enabling rapid adaptation to new or unseen tasks. Finn et al. [15] introduced a prominent meta-learning algorithm called Model-Agnostic Meta-Learning (MAML). MAML utilizes a fixed number of gradient descent steps on a meta-gradient, computed from various learning tasks (episodes), to achieve rapid parameter adaptation of the model for new tasks.
The concept of meta-learning can also be applied to graph adversarial attacks. As discussed in Section 2, attackers can leverage the meta-gradient obtained during training towards their target [16]. The meta-gradient is computed by unrolling the training process into multiple tasks, which can help the attack learn a more holistic picture of the model than looking into any certain phase in the training. By leveraging this meta-gradient, attackers gain a powerful road map, enabling them to manipulate graph data effectively and apply a powerful attack strategy. More specifically, the attacker can identify the node or links with the greatest impact during model learning and perform the precise attack instead of making random or naive perturbations on the graph. In this paper, we will take the advantage of the meta-learning to design our poisoning attack strategies.
Figure 2: Framework of the proposed approach.
4 Methodology
This section outlines the methodology of the proposed approach, including its objectives, attack settings, and attack framework.
4.1 Objective
In the proposed adversarial attack approach, the attacker utilizes meta-learning during the training phase of the attack model to make subtle modifications to the unweighted graph structure (adjacency matrix, denoted as $Adj$). As depicted in Fig. 2, these modifications generate poisoned data for the poisoning attack, which involves training the target model with the poisoned data—a process known as poisoning training.
The goal of this approach is to diminish the prediction accuracy of the model specifically on the targeted link(s). In other words, our objective is to reduce the link prediction accuracy of specified link(s) on Variational Graph Auto-Encoders through the poisoning attack.
4.2 Attack Setting and Threat Model
4.2.1 Attacker’s Goal
The attacker aims to increase the misclassification rate of the prediction of VGAE on a specific link (s) using the proposed poisoning attack. By introducing malicious and inconspicuous changes to the graph data, the attacker tries to make the learned VGAE model inherently biased after training on the poisoning data. For example, the goal is to induce the model to predict a link between two unrelated nodes, which should not be predicted.
4.2.2 Attacker’s Knowledge
In the original machine learning and deep learning process, there are typically three components: data, model, and training pipeline. In our approach, we assume that the attacker only has limited knowledge.
  • Graph Data Knowledge: The attacker has full access to the graph structure (adjacency matrix), node features, and node labels.
  • Model Knowledge: The attacker does not know the exact model architecture, such as the number of layers and hyperparameters of the model. They only know that the VGAE or GNN-based model is used for link prediction.
  • Training Pipeline Knowledge: The attack has a general understanding of when the model is trained or updated to time the poisoning attack. However, they do not have control or alter the original training procedure (i.e., the optimizer, learning rate, or objective function).
4.2.3 Attacker’s Capacity
In the proposed poisoning attack, the attacker has the ability to manipulate graph data with a limited budget:
  • Graph Structure Manipulation: The attacker can add or remove links in the adjacency matrix. The other modification to the graph data is not allowed.
  • Attack Budget: The attacker has a budget that restricts how many links can be added or removed. This budget typically scales with the size of the graph to allow sufficient modifications to be made on large graphs. By limiting the attack budget, the attacker can achieve inconspicuous modifications.
4.2.4 Threat Model
Leveraging the attacker’s knowledge and capacity, the attacker selectively adds or removes links within the allowed budget to generate poisoning data via their own framework (Section 4.3). These poisoned data distort the structural patterns of the graph, causing the VGAE to learn misleading latent representations. In other words, the attacker relies on subtle but carefully chosen structural modification on the graph data to achieve their goal of inducing incorrect link prediction.
4.3 Attack Framework
Fig. 2 provides an overview of our approach. It begins with the original graph data (original $Adj$ and original features) as input and generates poisoned graph data (poisoned $Adj$ and original features) for poisoning training. Our poisoning attack consists of two primary components that operate sequentially to produce the poisoned graph data for conducting poisoning training on VGAE:
  • Attack model: The attacker employs a surrogate model as an attack model to train on the original graph data. During the training, meta-learning with meta-gradient is utilized by the attacker to identify effective modifications to the graph structure.
  • Graph structure modifier: Following the training of the attack using meta-learning, the attacker employs a graph structure modifier to adjust the original adjacency matrix (original $Adj$) and create an adversarial adjacency matrix (adversarial $Adj$). Subsequently, the attacker combines the adversarial $Adj.$ with the original graph features to generate poisoned graph data for poisoning training.
Input: Attack budget $\Delta$, graph data $G(A,X)$, number of epoches $k$, attack model $M$
Output: Adversarial adjacency matrix $A_{\text{adv}}$
1 Initialize model parameters $\theta_{0}$ for $M$;
2
3for $i=0$ to $k$ do
4       Train $M$ and compute training loss $L$;
5       Compute the gradient $\nabla$ with respect to $\theta_{i}$ based on $L$;
6       Compute attack loss $\widetilde{L}_{i}$;
7       $\widetilde{L}$ = $\widetilde{L}$ + $\widetilde{L}_{i}$;
8       Back propagation on $\nabla$: update $\theta_{i}\rightarrow\theta_{i+1}$;
9
10Compute the meta-gradient $\widetilde{\nabla}$ with respect to attack target based on $\widetilde{L}$ ;
11 $A_{adv}=$ Adversarial modification(~$\widetilde{\nabla},A,\Delta$);
12
13return $A_{adv}$;
Algorithm 1 Adversarial Attack on VGAE using Attack Model with Meta-learning
Input: Gradient matrix $\widetilde{\nabla}$, adjacency matrix $A$, attack budget $\Delta$
Output: Adversarial adjacency matrix: $A_{\text{adv}}$
1 Initial $A_{adv}=A$
2 Symmetry the gradient matrix $\widetilde{\nabla}$ = ($\widetilde{\nabla}$ + ~$\widetilde{\nabla}^{T}$)/2
3 while $\Delta>0$ do
4       Get the index of the largest magnitude in $\widetilde{\nabla}$:$(i,j)$ = $|\widetilde{\nabla}|$);
5       if ~$\widetilde{\nabla}_{ij}>0$ and $A_{ij}=0$ then
6             $A_{adv}(i,j)=1$;
7             $\Delta$ = $\Delta$ 1;
8
9      if ~$\widetilde{\nabla}_{ij}<0$ and $A_{ij}=1$ then
10             $A_{adv}(i,j)=0$;
11             $\Delta$ = $\Delta$ 1;
12
13      Set position (i, j) immutable;
14
15return $A_{adv}$;
Algorithm 2 Adversarial Modification (Graph Structure Modifier)
4.4 Attack Strategy
Algorithms 1 and 2 outline the step-by-step process to generate the adversarial adjacency matrix through the attack model and make adversarial modifications using the graph structure modifier.
The primary purpose of the attack model is to extract information by mimicking the original process since the attacker lacks access to the original model and the training process. The essence of the attack model (Algorithm 1) lies in leveraging meta-learning during its training phase, enabling the attacker to gain a comprehensive understanding of the process. The attack model is trained on the original graph data using weighted cross-entropy (Algorithm 1, lines 3 and 4), defined as:
$L=\sum_{ij}-w{A_{ij}}\ln(\hat{A}_{ij})-(1-A_{ij})\ln(1-\hat{A}_{ij})$ (2)
where $w$ is the weight for weight cross-entropy, $w=(N^{2}-\sum_{ij}A_{ij})/\sum_{ij}$, $N$ represent the number of nodes, $A_{ij}$ is the label in original $Adj$, and $\hat{A}_{ij}$ is the predicted label.
The attacker minimizes this training loss with respect to the attack model parameter to optimize the performance of link prediction task. In addition, the attacker computes the attack loss in each training epoch (Algorithm 1, line 5), defined as:
$\tilde{L}=-w{Y}_{t}\ln(\tilde{A}_{t})-(1-{Y}_{t})\ln(1-\tilde{A}_{t})$ (3)
where $w$ is the weight for weight cross-entropy, $w=(N^{2}-\sum_{ij}A_{ij})/\sum_{ij}$, $N$ represent the number of nodes, ${Y}_{t}$ is the label in original $Adj$, and $\tilde{A}_{t}$ is the predicted label.
The attack loss is aggregated as meta-learning approach during each epoch (Algorithm 1, line 6). The key part of the attack model is that the attacker considers each training epoch as a task (episode) and intends to learn from all tasks (epoches). Instead of looking into a specific phase in the training process, meta-learning helps the attacker to capture the overall information of the whole training. The attacker tries to maximize the attack loss with respect to the graph structure ($Adj$) to conduct the effective attack on the graph structure. At the end of the training, the attacker computes the meta-gradient from the attack loss for further processing(Algorithm 1, line 8).
The meta-gradient obtained from the attack model training provides the attack direction for the graph structure modifier to apply the adversarial modification on the graph. Specifically, the values in the meta-gradient matrix represent their impact during model training, with larger magnitude (absolute values) indicating greater effectiveness. The largest magnitude of the meta-gradient points out the most effective link positions (Algorithm 2, line 4). Based on this, the attacker continues to modify the corresponding positions of the original $Adj.$ to acquire the adversarial $Adj.$ (Algorithm 2, lines 4-11) until the attack budget is exhausted. Since the graph structure in an unweighted graph is discrete, the graph structure modifier is limited to two actions: adding or removing links in the original $Adj.$ matrix. This involves altering the values of the adjacency matrix to either 0 (to remove a link) or 1 (to add a link). Furthermore, the attack budget imposes constraints on the total number of modifications allowed. By carefully regulating the quantity and nature of these modifications, the attack can execute unnoticeable poisoning attacks on the graph data that are hard to detect.
Table 1: Results of the Targeted Attack for Link Adversarial Dissolution.
Cora Citeseer Pubmed
Attack Budget
1% 2.5% 5% 1% 2.5% 5% 1% 2.5% 5%
Attack Approach ASR
Null Model [23] 10% 10% 10% 10% 10% 15% 10% 10% 15%
IGA [14] 25% 30% 30% 30% 30% 35% 25% 25% 30%
PGD [12] 35% 35% 40% 40% 40% 40% 35% 35% 40%
MetaBase [16] 40% 40% 45% 40% 40% 40% 35% 35% 45%
DICE [13] 30% 35% 40% 35% 35% 40% 35% 35% 35%
Our approach 45% 50% 50% 45% 45% 50% 40% 45% 50%
Table 2: Results of the Targeted Attack for Link Adversarial Activation.
Cora Citeseer Pubmed
Attack Budget
1% 2.5% 5% 1% 2.5% 5% 1% 2.5% 5%
Attack Approach ASR
Null Model [23] 15% 20% 20% 15% 15% 20% 15% 15% 15%
IGA [14] 55% 55% 60% 50% 60% 60% 45% 55% 60%
PGD [12] 60% 60% 60% 60% 60% 65% 55% 60% 60%
MetaBase [16] 65% 70% 70% 60% 65% 70% 65% 65% 65%
DICE [13] 60% 65% 65% 50% 60% 65% 55% 55% 60%
Our approach 100% 100% 100% 100% 100% 100% 90% 90% 90%
5 Experimental Results
In this section, we outline our experimental design and analyze the results obtained from our approach.
5.1 Experimental Setup
To evaluate the effectiveness and performance of our proposed attack method, we implemented the entire architecture using PyTorch. Our experiments aimed to answer the following questions:
  • Does our approach reduce the link prediction accuracy of VGAE?
  • Is our approach effective in various attack scenarios (objectives)?
  • How does our approach compare to the state-of-the-art baseline methods?
  • What is the impact of the attack budget on our approach?
To answer these questions, we deployed a two-layer VGAE to evaluate the effectiveness of our poisoning attack approach on three widely used graph datasets. By training VGAE separately on the original and poisoned graph data, we can observe any disparities in link prediction performance, indicating the effectiveness of our attack. In addition, to assess adaptability and robustness, we conducted evaluations under two distinct attack scenarios, each aligned with a specific attack objective. We compared the efficacy of our approach with six state-of-the-art methods, namely PGD [12], DICE [13], IGA [14], CLGA [17], MCGCL [24] and MetaBase [16]. Since CLGA [17] and MCGCL [24] are both designed based on the embedding of the graph that inherently considers the entire graph. It is challenging to isolate the single link as a target in their framework. Therefore, we limit our comparison of these two methods to untargeted attacks.
In our experiment, the attacker also utilizes a VGAE as the attack model to apply the adversarial modification on graph data and maximize the ability of our attack. As mentioned in the previous section, the adversarial modification is constrained by the attack budget, and we investigate the attack budget as a key parameter in our experiments. To ensure that the attack remains unnoticeable, we set the attack budget within a small range, varying from 1% to 5% of the total links in the dataset to produce the poisoned data for our poisoned attack. For each experiment, we conducted original training using original data and poisoning training using poisoned data. We then test both models with the same test data to evaluate their link prediction performance. The performance of the original model will be used as a baseline. The difference in performance between these two models indicates the effectiveness of our proposed attack. Moreover, we also leverage the null model [23] as an additional baseline. In the null model, the total number of links in the graph is fixed, and adversarial graphs are generated by randomly removing and reconnecting links within the original graph. Comparing our attack approach with the null model allows us to evaluate whether our approach explicitly highlights the significant vulnerable points in the graph structure.
Figure 3: Different adversarial attack scenarios under targeted attack.
5.2 Experimental Dataset
In our experiments, we utilized three widely used graph datasets sourced from public repositories to evaluate the performance of our proposed attack approach. These datasets represent scientific publications, each node representing a publication, and each link denoting citations.
  • Cora: The Cora dataset [25] contains scientific publications that are categorized into 7 classes. It has 2708 nodes and 5429 links. Each node is described by a word vector that indicates the presence of certain words in the publication from a predefined dictionary.
  • Citeseer: The Citeseer dataset [26] is a collection of scientific publications classified into 6 classes. It has 3312 nodes and 4732 links. Each node is described by a word vector that indicates the presence of certain words in the publication from a predefined dictionary.
  • Pubmed: The Pubmed dataset [27] comprises scientific publications from the PubMed database categorized into 3 classes. It has 19717 nodes and 44338 links. Each node is described by a Term Frequency-Inverse Document Frequency (TF-IDF) weighted word vector to capture the importance of words in the publication.
(a) Cora dataset
(b) Citeseer dataset
(c) Pubmed dataset
Figure 4: ROC-AUC Results of the Untargeted Attack Scenario.
(a) Cora dataset
(b) Citeseer dataset
(c) Pubmed dataset
Figure 5: AP Results of the Untargeted Attack Scenario.
5.3 Attack Scenarios
As elaborated in Section 3, adversarial attacks can manifest in two distinct scenarios (based on the objectives of the attack): targeted attacks and untargeted attacks. In our experiments, we address both scenarios as follows:
5.3.1 Targeted Attack
In this scenario, the attacker aims to manipulate the prediction of a single link within the graph, which can be categorized into two types as depicted in Fig. 3:
  • Link adversarial dissolution: In this attack, the objective is to alter/manipulate the prediction result of a link from present (1 - link exists) to absent (0 - link does not exist) in the adjacency matrix, thereby affecting the targeted position.
  • Link adversarial activation: In contrast, this attack seeks to change the prediction of a link from absent (0 - link does not exist) to present (1 - link exists) in the adjacency matrix.
Generally, when observing the performance of targeted attacks, most works focus on overall attack performance. However, it is crucial to investigate both link activation and dissolution caused by our attacks because these phenomena can have vastly different implications across various applications. For example, in social networks, the activation of links can foster stronger relationship ties, enhancing community engagement and collaboration. In contrast, the dissolution of links might isolate relationships, leading to fragmentation and decreased user interaction. Therefore, we evaluate both cases separately to help us better understand and address the potential vulnerabilities of link prediction.
5.3.2 Untargeted Attack
In this scenario, the attacker focuses on a specific set of links (e.g., testing set) of the entire graph, aiming to maximize the misclassification rate within that set.
5.4 Evaluation Metric
  • Attack Success Rate (ASR): For targeted attacks, ASR indicates the proportion of successful targeted attacks out of all attempted targeted attacks. This metric reflects the overall attack strength of the proposed attack, as shown in Eq. (4).
    $ASR=\frac{all\>success\>attack\>attampts}{all\>attack\>attampts}$ (4)
  • ROC-AUC and AP Score: In untargeted attacks, we utilize the Receiver Operating Characteristic (ROC), Area Under the Curve (AUC), and Average Precision (AP) Score to evaluate the performance of the model after the poisoning attack. ROC-AUC score provides a comprehensive understanding of the model’s ability to distinguish between positive and negative links. The ROC-AUC score is threshold-independent, evaluating model performance across all possible thresholds, and is more resilient to class imbalance compared to accuracy, making it a robust measure. Additionally, it provides a clear basis for comparing multiple models. Moreover, as graph data are usually considered imbalance data, the AP score emphasizes the precision-recall trade-off of the model. By comparing the original model’s performance with its post-attack performance, the effectiveness of the poisoning attack is clearly demonstrated, highlighting its impact on the robustness and performance of the model.
5.5 Effectiveness Analysis
5.5.1 Evaluation of targeted attack
Tables 1 and 2 present the success rates of 20 attack attempts. All the state-of-the-arts and our approach outperform the null model baseline. This shows that all these attack approaches identify and exploit vulnerable parts. A higher ASR indicates a stronger targeted attack ability on individual target links in the link prediction task. Our approach demonstrates remarkable effectiveness in the targeted attack scenario, outperforming other state-of-the-art methods. Additionally, our approach achieves the attacker’s objective with a smaller attack budget, making it more inconspicuous. As the attack budget increases in targeted attacks, there is a slight increment in the targeted attack results. This could be attributed to the fact that for a targeted attack on a single link, fewer modifications are needed to degrade the prediction accuracy of that link.
Another notable observation from the targeted attack results is that the ASR for link adversarial dissolution is significantly higher than that for link adversarial activation. This suggests that breaking an existing link is more challenging than creating a new adversarial link, both in our attack and in other state-of-the-art methods. This intuitively aligns with real-world scenarios; forming a new connection between two entities is generally simpler than breaking an existing relationship. For instance, in a social network, it is generally simpler to establish a new connection between two individuals than to dissolve an existing relationship between them.
5.5.2 Evaluation of untargeted attack
The results of the untargeted attack in the three datasets are depicted in Fig. 4 and Fig. 5. In the experiments, we used two baselines, the original model and the null model. The original model represents the performance of the link prediction on the original graph data. It is unaffected by the attack budget as it is clean data. The effectiveness of the attack methods is evaluated by observing the drop in ROC-AUC and AP scores. The null model serves as a benchmark to evaluate whether the approach is capable of highlighting the vulnerabilities of the graph data. All approaches demonstrate their effectiveness compared to the original model and null model. Our method shows superior performance in untargeted attack scenarios, slightly surpassing other state-of-the-art approaches. In particular, our method achieves a lower ROC-AUC and AP score while requiring a smaller attack budget. Unlike targeted attacks, the impact of the attack budget is more pronounced in untargeted attacks. As the attack budget increases, the ROC-AUC and AP score experiences a significant drop.
(a) Original Graph
(b) Poisoned Graph (5%)
(c) Poisoned Graph (10%)
(d) Poisoned Graph (20%)
Figure 6: Visualization Examples of original graph and poisoned graph resulting from the proposed attack.
Graph Statistic Original Graph Poisoned Graph (5%) Poisoned Graph (10%) Poisoned Graph (20%)
Number of Nodes 2708 2708 2708 2708
Number of Edges 4488 4656 4986 5358
Avg. Degree 3.42 3.56 3.80 4.15
Graph Density 0.001 0.001 0.001 0.002
Network Diameter 21 21 18 14
Avg. Clustering Coefficient 0.24 0.25 0.27 0.30
Avg. Path Length 6.79 6.42 5.80 5.00
Table 3: Comparison of statistics of the two graphs.
5.6 Graph Visualization
Fig. 6 shows a visualization example of the original graph and the poisoned graph under different attack budgets (ratio of the total number of edges) resulting from the proposed attack using the Cora dataset. All visualizations were generated using Gephi with the same layout setting and fixed node coordinates. To better present the local structure of the graph data, we also applied the community detection results on the visualization example. The nodes are colored according to their community labels, with the labels ranked by community size from largest to smallest as follows: purple, green, blue, black, orange, and red. The purpose of visualizing these graphs is to investigate how the proposed attack affects the original graph. In other words, we do not expect the proposed attack to produce conspicuous modifications to the original graph. As we mentioned in the previous section, our approach limits the attack budget within 5% of the total links in the graph to achieve the unnoticeable changes in the original graph. The visualizations demonstrate that the attacker applies the minor changes on the original graph, which can be considered as unnoticeable.
Upon observing Fig. 6 (a) and 6 (b), it is apparent that the two graph visualizations look very similar. Most of the communities are almost identical, indicating that the local structure of the graph is preserved. Since the ratio of modified links is much smaller than the total number of links, it is challenging to discern the detailed adversarial link modifications resulting from the proposed poisoned attack. These changes can be considered unnoticeable in the original graph. However, when observing Fig. 6 (c) and 6 (d), we notice that the overall appearance appears somewhat different. Most of the communities have changed, and even the sizes of the communities have changed. Upon closer inspection, Figs. 6 (c) and 6 (d) are more interconnected than Figs. 6 (a) and 6 (b). In addition, some new links have appeared in the upper right corner. To further investigate, we calculated the statistics of these graphs, as shown in Table 3.
From Table 3, we observe that the statistics of Figs. 6 (a) and 6 (b) are very similar, only have slight differences in some aspects. The statistic of Fig. 6 (c) and 6 (d) have much more difference, especially on Fig. 6 (d). We also conclude some critical findings from the visualization and statistic table. Firstly, we noticed that our adversarial modifications add more links to the original graph. Secondly, the average degree, average clustering coefficient, and average path length also indicate that the poisoned graph is more interconnected than the original graph. As the attack budget increases, the graph becomes more interconnected. These observations collectively demonstrate that the proposed attack slightly enhances the connectivity of the original graph. This enhancement is likely attributable to the fact that link adversarial activation performs better than link adversarial dissolution in targeted attacks.
6 Conclusion and Future work
In this paper, we introduce an innovative unweighted graph poisoning attack leveraging meta-learning techniques to undermine the link prediction performance of VGAE. Our approach capitalizes on meta-learning to identify optimal attack directions, ensuring subtle modifications in the graph. Through extensive experimental evaluation, we showcased the efficacy of our attack in targeted and untargeted scenarios. Our results demonstrate superiority over existing state-of-the-art methods.
For future research, we plan to explore three potential directions:
  • Improving link adversarial dissolution: Our approach exhibits limited effectiveness in breaking existing links. Future work could focus on improving the performance of this aspect of the attack.
  • Application in recommendation systems: Investigate the applicability of our approach in real-world recommendation systems, which often rely on user-item graphs for personalized recommendations.
  • Extension to weighted graph attack: Extending our current work to address graph adversarial attacks on weighted graphs expands the scope of our approach to handle more complex graph structures.